OpenStudy (baby456):

Please help Me Asap! Medal + Fan! A farmer plants apple trees. The total number of apples can be represented by the function A(t)=300t-5t^2. Where t is the number of trees he plants. a. Sketch the graph of A(t)=300t-5t^2 on the coordinate plane belw. Show the points where t= 0, 10,20,30, 40, 50,60. b. The farmer wants to be able to harvest at least 4000 apples. Solve the inequality 300t-5t^2=>4000.

OpenStudy (baby456):

@texaschic101

OpenStudy (baby456):

@jim_thompson5910 this is a really hard question!

OpenStudy (baby456):

@jdoe0001 hard question!

OpenStudy (baby456):

This question is super @satellite73

OpenStudy (baby456):

super hard question isn't @lilcece

OpenStudy (baby456):

Could you help me @satellite73

OpenStudy (baby456):

algebra teacher gives us treally hard probelms!

OpenStudy (anonymous):

it is a parabola that opens down http://www.wolframalpha.com/input/?i=300t-5t^2+domain+0..60

OpenStudy (anonymous):

at \(t=0\) you have \(A(t)=0\) if you plant no trees, you get no apples

OpenStudy (anonymous):

at \(t=10\) you get \[A(10)=300\times 10-5\times 10^2=2500\] apples

OpenStudy (anonymous):

at \(t=20\) you get \[A(20)=300\times 20-5\times 20^2=4000\] apples

OpenStudy (anonymous):

etc

OpenStudy (baby456):

thank you so much!

OpenStudy (anonymous):

at \(t=40\) you get \[A(40)=300\times 40-5\times 40^2=4000\] apples as well that helps you answer the last question

OpenStudy (baby456):

how do i solve the inequality!

OpenStudy (anonymous):

The farmer wants to be able to harvest at least 4000 apples. Solve the inequality \(300t-5t^2\geq4000\)

OpenStudy (anonymous):

you see that you get \(4000\) at both \(t=20\) and also at \(t=40\) more in between (from the picture) so the solution is \(20\leq t\leq 40\) i.e. between 20 and 40 trees

OpenStudy (baby456):

i get it so much better now thank you so much!

OpenStudy (anonymous):

yea its hard im only in the 6th grade